OpenStudy (anonymous):

Please help! What is the value of the y variable in the solution to the following system of equations? 5x + 4y = 1 4x + 3y = -1 Choices: -7 7 9 -9

OpenStudy (nurali):

5x + 4y = 1---------(1) 4x + 3y = -1--------(2) in equation (1) 5x + 4y = 1 5x=1-4y x=1-4y/5--------(3) put 'x' in equation (2) 4x + 3y = -1 4(1-4y/5)+3y=-1 4-16y/5+3y=-1 4-16y+15y=-5 4-y=-5 y=9

OpenStudy (anonymous):

Thanks!